6/X + .14 = .29 Please help me with this thanks so much!
Answer by checkley77(12569)

6/X + .14 = .29
6/X=.29-.14
6/X=.15
.15X=6
X=6/.15
X=40 ANS.
PROOF:
6/40+.14=.29
.15+.14=.29
.29=.29